OpenStudy (anonymous):

Tony has an order for a rectangular sheet of metal the measures 42m^2. If he already has a sheet that is 7 meters long and 2 meters wide what size sheet will he need to weld to this sheet in order to complete the order?

OpenStudy (anonymous):

Area of sheet to send = 42 m^2 Area of sheet present = 7*2 m^2 = 14 m^2 Area more required = 42-14 = 28 m^2